One of the performance characteristics of a car is the time needed to accelerate from 0 km·h-1

to 100 km·h-1. Consider a car that can do it in 4.0 s. Find the distance travelled by the car

while accelerating. Suppose that the acceleration of the car is constant.


1
Expert's answer
2021-02-02T19:46:32-0500

We can find the distance traveled by the car while it accelerating from the kinematic equation:


"d=\dfrac{1}{2}(v_i+v_f)t,""d=\dfrac{1}{2}(0+100\ km\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s})\cdot 4\ s=55.5\ m."